Klapperich Lab:Notebook/Lab Meeting Notes/2010/04/22

From OpenWetWare
Jump to navigationJump to search
Owwnotebook icon.png Project name Report.pngMain project page
Resultset previous.pngPrevious entry      Next entryResultset next.png

22 April 2010 Lab Meeting

‡ Announcements

  • May 6 Lab meeting cancelled
  • Send me any conflicts for summer lab meeting time.
  • Sr. Project Presentation practice at 4/29 meeting.

‡ Flu R01:Integration

  • meetings every other M 1-3pm. 705.
  • QQ gone at Oakridge conf.
  • for PCR just do more reactions, pool and etoh precipitate.
    • For HDA, we need to explore cloning. Waiting for the competent cells.
    • dxs PCR primers worked for HDA in tube. Still no positive results on chip.
  • Check other methods of heating.

† HDA (Lead:Sonali)

  • IBC Approval or 720 to work with real C.diff samples.
  • Patient Samples - dilution experiments (MM/SH)
  • Start planning R01 for Submission on 10/5/10. MM, CMK.
  • HDA mix in dried form - Expecting it from BH -- ON HOLD.

  • QQ preliminary chip work
  • Sequence all output
  • 2nd HDA design - "microSPE" + "Flu primer" + "HDA in a cup" + "Detection kit from Biohelix"

† Sample Concentration (Lead: Jane, Team: Jaephil)

  • Straight channel adhesion test PCR result expected. From 4/6 result, PDMS 4mm wide channel doesn't show decrease in output RNA over time
  • Straight 1mm and 4mm PDMS-Done
  • Try collection vs dilutions, collection vs time, need binding constant to quantify results.

† SPE Column Optimization for DNA/RNA. (Lead: )

  • (Almost done) Jessie will run chips with serial virus dilutions from 1:1 to 1:1,000,000 to re-do previous dilution experimnent with WHO SYBR green assay on chip.

† Integrated chip for flu(SPE+RT+PCR) (Lead: Qingqing)

  • Real copy number of previous human samples. So QQ is working with the higher concentration samples. Range is 10^6-10^8 of copy number/ml.
  • REVISIT the % recovery number.
  • QQ tried: more enzyme, Tween 20, BSA amounts.
  • integrated flu chip paper has been send to cathie, wait for revision.

  • PCR1 will combine with PCR2, waitting for revision (CMK)

‡C. diff Project (Cathie, Sonali, Satish Singh, Lisa J., His post doc )

  • See email notes.

‡ Coulter Flu Fraunhofer Project (Lead: Sonali, Team: Sonali, Jessie, Cathie, CMI Folks, Qingqing)

  • New meeting time, T, 9:30 am every other week. MEETINGS CANCELLED NEXT WEEK (4/20).

‡ Agilent Automated Sample Preparation (Lead: Alex)

  • Paper on HotDog - with CMK/ASB

‡ COBRA (Lead: Jaephil, Team: Cathie, Jane for virus only, PHO folks)

  • Evap paper - Additional experiment for Fig 3. JD says will send by Friday (4/16).

‡ Biointerfaces group (Lead: MinCheol, Team: Cathie, MCK, Wong and Meller folks)
* Cathie will submit paper.

‡ CIMIT- Sepsis (Lead:Cathie, Team:TBA)

  • Sample collection ongoing.

‡ PATH Grant (Lead:Cathie, Team:Frank, Sean, Mark, Jake Trueb (ME))

  • .
  • Frank: Four-week storage results sent to PATH group. +/- carrier RNA results sent as well -- see increased ng amount extracted withOUT RNA, but not statistically significant (~60 vs. 40ng). Longer-term (2, 3, 4-month) in storage. Currently writing up final report for senior project. After classes end, will be compiling documents/organizing supplies for whoever is taking over the project.
  • Sean: Data for 700 nm and 150 nm analyzed and it shows that 700 is better overall, and is best in the 60 - 80 SPE range. Final report almost complete and due tomorrow (4/23). I will get stuff together and leave my notebook for future research as well.
  • Mark: as of 4/15/10 up to 30% recovery from 10 ng input sample with water elution. Also up to 13% recovery with 100ng/straw. 1-week trial completed as of 3/24. Results peak at 20% (-70C) and decrease with temperature. 2-month and 1-month trials complete, resulting in further degradation than in the 1-week trial. 4- and 6-month trials still in storage. Final Report draft submitted, due tomorrow. Presentation next Friday 4/30. Lab notebook will be updated and ready for future PATH SNAP workers.

‡ IIH Senior Project.

  • Ordered new HRP substrate. waiting for delivery
  • Will run ELISA in chip and image for standard curve.
  • finalizing report